### v1

# Y. Grandati1 and C. Quesne2

Abstract
A previous construction of regular rational extensions of the trigonometric Darboux-Pöschl-Teller potential, obtained by one-step Darboux transformations using seed functions associated with the para-Jacobi polynomials of Calogero and Yi, is generalized by considering $m$-step Darboux transformations. As a result, some novel families of exceptional orthogonal polynomials depending on $m$ discrete parameters, as well as $m$ continuous real ones $\lambda_{1}$, $\lambda_{2}$, …, $\lambda_{m}$, are obtained. The restrictions imposed on these parameters by the rational extensions regularity conditions are studied in detail.
I INTRODUCTION
Exceptional orthogonal polynomials (EOPs) are complete families of orthogonal polynomials that arise as eigenfunctions of a Sturm-Liouville eigenvalue problem [1]. Their most apparent difference with respect to the classical orthogonal polynomials (COPs) of Jacobi, Laguerre, and Hermite is that there are some gaps in the sequence of their degrees, the total number of missing “exceptional” degrees being known as the codimension. In addition, the corresponding differential equation contains rational coefficients instead of polynomial ones.
Since their appearance, EOPs have generated a lot of research activity both from a mathematical viewpoint and for their applications in mathematical physics. In mathematics, studies have been carried out on the properties of their zeros [2, 3], the sets of recurrence relations they satisfy [4, 5, 6], and the search for a full classification [7, 8]. In mathematical physics, EOPs were shown to be related to the Darboux transformation (DT) in the construction of exactly solvable rationally-extended quantum potentials [9, 10, 11]. This led to multi-indexed families of EOPs connected with multi-step DT [12, 13, 14, 15]. EOPs also appeared in connection with exact solutions of the Dirac equation [16], some superintegrable systems [17, 18], or rational solutions of some Painlevé equations [19].
Before 2015, it was thought that apart from the parameters associated with the corresponding COPs, there only appeared discrete parameters in the construction of EOPs. In the Jacobi family case, for instance, the most general known construction was given in terms of a Wronskian-like determinant of classical Jacobi polynomials, indexed by two partitions [20, 21]. Some possibilities of having more general EOP families, also depending on continuous parameters, then made their appearance.
First, on considering the trigonometric Darboux-Pöschl-Teller (TDPT) potential, whose bound states are expressed in terms of Jacobi polynomials, and using the fact, first noticed by Szegö [22] and emphasized by Calogero and Yi [23], that for some values of the parameters the Jacobi equation has general polynomial solutions depending on an arbitrary real parameter $\lambda$ and called para-Jacobi polynomials, it was possible to build some new regular rational extensions of the TDPT potential by one-step DT [24]. The eigenstates of such extensions were shown to be associated with a novel family of $\lambda$-dependent EOPs.
Second, in the cases of the TDPT and isotonic potentials, some new rational extensions were constructed by considering two-step confluent chains of DT, that is chains of DT in which the spectral parameters of the different seed functions converge to the same value [25]. Both types of extended potentials are associated with new families of orthogonal polynomials. In the case of the TDPT potential, the latter depend on a continuous real parameter. Later on, the confluent DT algorithm was also used to construct exceptional Legendre polynomials [26], as well as exceptional Gegenbauer ones [27] with an arbitrary number of real parameters.
Third, a dualization of Krall dual Hahn polynomials was used to construct exceptional Hahn polynomials, which in the limit led to exceptional Jacobi polynomials depending on an arbitrary number of continuous parameters [28]. The latter were then shown [29] to be divided in two classes: some families that are deformations of Jacobi polynomials, as those determined in [25], and other families that are deformations of standard families of Jacobi EOPs.
The purpose of the present paper is to complete the study made in Ref. [24], which was limited to one-step DT. Here, we plan to study in detail its multi-step version, which will allow us to derive novel families of exceptional para-Jacobi polynomials depending on an arbitrary number of free continuous parameters.
The paper is organized as follows. We start by recalling the essential features of DT and of the TDPT potential in Secs. II and III, respectively. In Sec. IV, we summarize the results obtained for one-step regular extensions of the TDPT potential based on para-Jacobi polynomials. In Sec. V, we study in detail two-step regular extensions of the same. In Sec. VI, a similar analysis is carried out for $m$-step regular extensions. Finally, Sec. VII contains the conclusion.
II DARBOUX TRANSFORMATIONS (DT)
A One-step DT
We consider a one-dimensional Hamiltonian $\widehat{H}=-d^{2}/dx^{2}+V(x)$, $x\in I\subset\mathbb{R}$ and the associated Schrödinger equation
$\psi_{\lambda}^{\prime\prime}(x)+(E_{\lambda}-V(x))\psi_{\lambda}(x)=0,$ (2.1)
where $\psi_{\lambda}(x)$ is a formal eigenfunction of $\widehat{H}$ for the eigenvalue $E_{\lambda}$ and is defined up to a multiplicative constant. In the following, we suppose that with Dirichlet boundary conditions on $I$, $\widehat{H}$ admits a discrete spectrum of energies and eigenstates $(E_{n};\psi_{n})_{n\in\left\{0,...,n_{\max}\right\}\subset\mathbb{N}}$ where, without loss of generality, we can always suppose that the ground level of $\widehat{H}$ is zero: $E_{0}=0$.
Any formal eigenfunction $\psi_{\nu}(x)$ of $V(x)$(i.e., of $\widehat{H}$) can be used as a seed function for a Darboux transformation (DT) $A(\psi_{\nu})$, which associates with the potential $V(x)$ a modified potential
$V(x)\overset{A(\psi_{\nu})}{\longrightarrow}V^{(\nu)}(x)=V(x)-2(\log(\psi_{\nu}(x)))^{\prime\prime},$ (2.2)
which we call an extension of $V(x)$. The formal eigenfunction of $V^{(\nu)}$ associated with the spectral parameter $E_{\lambda}$ is given by the Darboux-Crum formulas [30]
$\left\{\begin{array}[]{c}\psi_{\lambda}^{\left(\nu\right)}(x)\sim\frac{W\left(\psi_{\nu},\psi_{\lambda}\mid x\right)}{\psi_{\nu}(x)},\text{ if }\lambda\neq\nu,\\ \psi_{\nu}^{\left(\nu\right)}(x)\sim\frac{1}{\psi_{\nu}(x)},\end{array}\right.$ (2.3)
where $W(y_{1},...,y_{m}\mid x)$ denotes the Wronskian of the family of functions $y_{1}(x),...,y_{m}(x)$ [31].
$W\left(y_{1},...,y_{m}\mid x\right)=\left|\begin{array}[]{ccc}y_{1}\left(x\right)&...&y_{m}\left(x\right)\\ ...&&...\\ y_{1}^{\left(m-1\right)}\left(x\right)&...&y_{m}^{\left(m-1\right)}\left(x\right)\end{array}\right|.$ (2.4)
This last possesses the following useful properties [31]
$\left\{\begin{array}[]{c}W\left(uy_{1},...,uy_{m}\mid x\right)=u(x)^{m}W\left(y_{1},...,y_{m}\mid x\right)\\ W\left(y_{1},...,y_{m}\mid x\right)=\left(\frac{dz}{dx}\right)^{m(m-1)/2}W\left(y_{1},...,y_{m}\mid z\right).\end{array}\right.$ (2.5)
B Chains of DT
At the formal level, the DT can be straightforwardly iterated and a chain of $m$ DT can be completely characterized by the $m$-tuple $(\nu_{1},...,\nu_{m})$ of spectral indices of the successive seed functions used in the chain. We denote such a $m$-tuple by a capital letter $N_{m}$, where the index $m$ indicates the length of the chain $N_{m}$. Let $\psi_{\lambda}^{\left(N_{m}\right)}$ be the formal eigenfunction associated with the eigenvalue $E_{\lambda}$ of the potential $V^{\left(N_{m}\right)}(x)$.
A chain is non-degenerate if all the spectral indices $\nu_{i}$ of the chain $N_{m}$ are distinct and is degenerate if some of them are repeated in the chain. For non-degenerate chains, Crum has derived very useful formulas for the extended potentials and their eigenfunctions in terms of Wronskians of eigenfunctions of the initial potential [30].
Crum’s formulas
When all the $\nu_{j}$ and $\lambda$ are distinct, we have
$\psi_{\lambda}^{\left(N_{m}\right)}(x)=\frac{W^{\left(N_{m},\lambda\right)}\left(x\right)}{W^{\left(N_{m}\right)}\left(x\right)}$ (2.6)
and
$V^{\left(N_{m}\right)}(x)=V(x)-2\left(\log W^{\left(N_{m}\right)}\left(x\right)\right)^{\prime\prime},$ (2.7)
where $W^{\left(N_{m}\right)}\left(x\right)=W\left(\psi_{\nu_{1}},...,\psi_{\nu_{m}}\mid x\right)$.
These formulas can be extended to degenerate chains if we adopt the convention to suppress any pair of repeated indices in the set of spectral indices associated with the chain.
III TDPT POTENTIAL
A Spectrum
The trigonometric Darboux-Pöschl-Teller (TDPT) potential (with zero ground-state energy) is defined on $]0,\pi/2[$ by
$V(x;\alpha,\beta)=\frac{\left(\alpha+1/2\right)(\alpha-1/2)}{\sin^{2}x}+\frac{\left(\beta+1/2\right)(\beta-1/2)}{\cos^{2}x}-(\alpha+\beta+1)^{2},$ (3.1)
and is a confining potential for $\left|\alpha\right|,\left|\beta\right|>1/2$. On introducing the variable $z=\cos 2x$, it can be rewritten as
$V(x;\alpha,\beta)=\frac{2\left(\alpha+1/2\right)(\alpha-1/2)}{1-z}+\frac{2\left(\beta+1/2\right)(\beta-1/2)}{1+z}-(\alpha+\beta+1)^{2}.$ (3.2)
For $\alpha,\beta>1/2$, the physical spectrum of the TDPT potential associated with the asymptotic Dirichlet boundary conditions
$\psi\left(0^{+};\alpha,\beta\right)=0=\psi\left(\left(\pi/2\right)^{-};\alpha,\beta\right)$ (3.3)
is given in terms of Jacobi polynomials $P_{n}^{\left(\alpha,\beta\right)}(z)$ by $\left(\left(\alpha_{n},\beta_{n}\right)=\left(\alpha+n,\beta+n\right)\right)$
$\left\{\begin{array}[]{c}\psi_{n}\left(x;\alpha,\beta\right)=\psi_{0}\left(x;\alpha,\beta\right)P_{n}^{\left(\alpha,\beta\right)}\left(z\right)\\ E_{n}\left(\alpha,\beta\right)=\left(\alpha_{n}+\beta_{n}+1\right)^{2}-\left(\alpha+\beta+1\right)^{2}=4n\left(\alpha+\beta+1+n\right)\end{array}\right.,n\in\mathbb{N},$ (3.4)
where
$\psi_{0}\left(x;\alpha,\beta\right)=2^{\alpha+\beta+1}\left(\sin x\right)^{\alpha+1/2}\left(\cos x\right)^{\beta+1/2}=(1-z)^{\left(\alpha+1/2\right)/2}(1+z)^{\left(\beta+1/2\right)/2}$ (3.5)
and
$P_{n}^{\left(\alpha,\beta\right)}\left(z\right)$ $=$ $\frac{1}{2^{n}}\sum_{k=0}^{n}\left(-1\right)^{n-k}\binom{n+\alpha}{k}\binom{n+\beta}{n-k}(1-z)^{n-k}(1+z)^{k}$ (3.6)
$=$ $\frac{\left(-1\right)^{n}\Gamma\left(n+\beta+1\right)}{n!\Gamma\left(n+\alpha+\beta+1\right)}\sum_{k=0}^{n}\left(-1\right)^{k}\binom{n}{k}\frac{\Gamma\left(n+\alpha+\beta+1+k\right)}{2^{k}\Gamma\left(\beta+1+k\right)}(1+z)^{k}.$
The quasipolynomial eigenfunction (i.e., which is, up to a gauge factor, polynomial in the adapted variable $z$) $\psi_{n}\left(x;\alpha,\beta\right)$ is then the particular solution of the Schrödinger equation
$\left(-\frac{d^{2}}{dx^{2}}+V\left(x;\alpha,\beta\right)-E_{n}\left(\alpha,\beta\right)\right)\psi\left(x\right)=0,$ (3.7)
which satisfies the Dirichlet boundary conditions (3.3).
B Para-Jacobi polynomials
Suppose that $\alpha$ and $\beta$ are two positive integers
$\alpha=N\in\mathbb{N}^{\ast},\qquad\beta=M\in\mathbb{N}^{\ast}.$ (3.8)
In this case, as noticed by Szegö [22] and emphasized by Calogero and Yi [23, 24], for values of $n$ such that [32]
$\max\left(N,M\right)\leq n<N+M,$ (3.9)
one obtains the general solution of the Schrödinger equation
$\left(-\frac{d^{2}}{dx^{2}}+V\left(x;N,M\right)-E_{-n-1}\left(N,M\right)\right)\psi\left(x\right)=0,$ (3.10)
which is of the quasipolynomial form
$\psi_{-n-1}\left(z;N,M;\lambda\right)=\psi_{-1}\left(z;N,M\right)p_{n}^{\left(-N,-M\right)}\left(z;\lambda\right),$ (3.11)
where, up to a constant factor,
$\psi_{-1}\left(z;N,M\right)$ $=\psi_{0}\left(z;-N,-M\right)$
$=(1-z)^{\left(-N+1/2\right)/2}(1+z)^{\left(-M+1/2\right)/2}$
$=\frac{1}{\psi_{0}\left(z;N-1,M-1\right)}.$ (3.12)
In Eq. (3.11), $p_{n}^{\left(-N,-M\right)}\left(z;\lambda\right)$ is a polynomial depending on an arbitrary real parameter $\lambda$ and is called para-Jacobi polynomial (In the following, $\psi_{-n-1}$ will be called a para-Jacobi (PJ) eigenfunction). It has the (monic) form [23]
$p_{n}^{\left(-N,-M\right)}\left(z;\lambda\right)$ $=$ $\frac{\left(-2\right)^{n}\left(n-M\right)!n!}{\left(2n-M-N\right)!}\Theta_{n,1}^{\left(-N,-M\right)}(z)$ (3.13)
$+\lambda\frac{\left(-2\right)^{n}\left(2n-M-N+1\right)!\left(M+N-n-1\right)!}{\left(n-N\right)!}\Theta_{n,2}^{\left(-N,-M\right)}(z),$
where
$\left\{\begin{array}[]{c}\Theta_{n,1}^{\left(-N,-M\right)}(z)=\sum_{k=M}^{n}\frac{\left(-1\right)^{k}\left(n-M-N+k\right)!}{2^{k}k!\left(k-M\right)!\left(n-k\right)!}(1+z)^{k}\\ \Theta_{n,2}^{\left(-N,-M\right)}(z)=\sum_{k=0}^{N+M-n-1}\frac{\left(-1\right)^{k}\left(M-1-k\right)!}{2^{k}k!\left(N+M-n-1-k\right)!\left(n-k\right)!}(1+z)^{k}\end{array}\right.,$ (3.14)
with $N,M>0$, and
$\left\{\begin{array}[]{c}\Theta_{n,1}^{\left(-N,-M\right)}(-1)=0\\ \Theta_{n,2}^{\left(-N,-M\right)}(-1)=\frac{\left(M-1\right)!}{\left(N+M-n-1\right)!n!}\end{array}\right..$ (3.15)
We have in particular
$p_{n}^{\left(-N,-M\right)}\left(-1;\lambda\right)=\left(-1\right)^{n}\lambda b_{n}^{\left(N,M\right)},$ (3.16)
where
$b_{n}^{\left(N,M\right)}=\frac{2^{n}\left(2n-N-M+1\right)!\left(M-1\right)!}{n!\left(n-N\right)!}.$ (3.17)
The para-Jacobi polynomials satisfy the derivation property
$\overset{.}{p}_{n}^{\left(-N,-M\right)}\left(z;\lambda\right)=np_{n-1}^{\left(-N+1,-M+1\right)}\left(z;a_{n}^{\left(N,M\right)}\lambda\right),$ (3.18)
with
$a_{n}^{\left(N,M\right)}=\frac{M+N-n-1}{n}.$ (3.19)
We also have the symmetry property [23]
$p_{n}^{\left(-N,-M\right)}\left(-z;\lambda\right)=\left(-1\right)^{n}p_{n}^{\left(-M,-N\right)}\left(z;g_{n}^{\left(N,M\right)}\left(\lambda\right)\right),$ (3.20)
where $g_{n}^{\left(N,M\right)}$ is the affine function
$g_{n}^{\left(N,M\right)}\left(\lambda\right)=\left(-1\right)^{n-M}\left(\left(-1\right)^{n-N+1}\lambda+\lambda_{n}^{\left(N,M\right)}\right),$ (3.21)
with
$\lambda_{n}^{\left(N,M\right)}=\frac{n!\left(n-M\right)!\left(n-N\right)!}{\left(2n-N-M\right)!\left(2n-N-M+1\right)!\left(N+M-n-1\right)!}>0.$ (3.22)
This implies in particular (cf Eq. (3.16))
$p_{n}^{\left(-N,-M\right)}\left(1;\lambda\right)=b_{n}^{\left(N,M\right)}g_{n}^{\left(N,M\right)}\left(\lambda\right).$ (3.23)
Note the following useful identities
$\lambda_{n-1}^{\left(N-1,M-1\right)}=a_{n}^{\left(N,M\right)}\lambda_{n}^{\left(N,M\right)},$ (3.24)
$b_{n-1}^{\left(N-1,M-1\right)}=\frac{n}{2\left(M-1\right)}b_{n}^{\left(N,M\right)},$ (3.25)
and
$b_{n}^{\left(M,N\right)}=\frac{\left(N-1\right)!\left(n-N\right)!}{\left(M-1\right)!\left(n-M\right)!}b_{n}^{\left(N,M\right)}.$ (3.26)
IV ONE-STEP REGULAR EXTENSIONS OF THE TDPT POTENTIAL
For $\max(N,M)\leq n<N+M$, we have $E_{-n-1}(N,M)<0$ and $\psi_{-n-1}\left(z;N,M;\lambda\right)$ is disconjugated on $x\in$ $]0,\pi/2[$ $(z\in$ ]1,+1[)]-1,+1[)] - 1 , + 1 [ ) and consequently admits at most one zero on this interval. As proven in [24], $\psi_{-n-1}$ has no node there iff the parameters satisfy the following regularity conditions:
$\left\{\begin{array}[]{c}(M,n-N)\in\left(2\mathbb{N}\right)^{2}:0<\lambda<\lambda_{n}^{\left(N,M\right)};\\[5.69046pt] (M,n-N)\in 2\mathbb{N}\times\left(2\mathbb{N}+1\right):\lambda<-\lambda_{n}^{\left(N,M\right)}\text{ or }\lambda>0;\\[5.69046pt] (M,n-N)\in\left(2\mathbb{N}+1\right)\times 2\mathbb{N}:\lambda<0\text{ or }\lambda>\lambda_{n}^{\left(N,M\right)};\\[5.69046pt] (M,n-N)\in\left(2\mathbb{N}+1\right)^{2}:-\lambda_{n}^{\left(N,M\right)}<\lambda<0.\end{array}\right.$ (4.1)
With the choices above, we can then use $\psi_{-n-1}$ as seed function of a one-step DT to produce a rational extension of the TDPT potential, which is perfectly regular on $]0,\pi/2[$. This extended potential is given by [24]
$V^{(-n-1)}\left(x;N,M;\lambda\right)=V\left(x;N-1,M-1\right)-4(N+M)-2\frac{d^{2}}{dx^{2}}\log(p_{n}^{\left(-N,-M\right)}\left(z;\lambda\right))$ (4.2)
and can be continuously modulated by varying the parameter $\lambda$ on the adapted interval.
Its spectrum is given by $E_{k}\left(N,M\right)$, $k\in\left\{-n-1,0,1,...,\right\}$, with the corresponding eigenstates:
$\psi_{-n-1}^{\left(-n-1\right)}\left(x;N,M;\lambda\right)=\frac{1}{\psi_{-n-1}\left(x;N,M;\lambda\right)}=\psi_{0}\left(x;N-1,M-1\right)\frac{1}{p_{n}^{(-N,-M)}(z;\lambda)}$ (4.3)
and
$\psi_{k}^{\left(-n-1\right)}\left(x;N,M;\lambda\right)$
$=\frac{W\left(\psi_{-n-1}(x;N,M;\lambda),\psi_{k}(x;N,M)\mid x\right)}{\psi_{-n-1}\left(x;N,M;\lambda\right)}$
$=\psi_{-1}(x;N,M)\frac{dz}{dx}\frac{W\left(p_{n}^{(-N,-M)}(z;\lambda),(1-z)^{N}(1+z)^{M}P_{k}^{(N,M)}(z)\mid x\right)}{p_{n}^{(-N,-M)}(z;\lambda)}$
$\ \propto\frac{\psi_{-1}(x;N-1,M-1)}{p_{n}^{(-N,-M)}(z;\lambda)}$
×$\times\left|\begin{array}[]{cc}p_{n}^{(-N,-M)}(z;\lambda)&(1-z)^{N}(1+z)^{M}P_{k}^{(N,M)}(z)\\ np_{n-1}^{\left(-N+1,-M+1\right)}\left(z;a_{n}^{(N,M)}\lambda\right)&-2(k+1)(1-z)^{N-1}(1+z)^{M-1}P_{k+1}^{(N-1,M-1)}(z)\end{array}\right|$ (4.6)
for $k=0,1,\ldots$. Here we have used Eq. (3.18), Eq. (2.5), and the following derivation rule of Jacobi polynomials [34]
$\frac{d}{dz}\left((1-z)^{N}(1+z)^{M}P_{k}^{(N,M)}(z)\right)=-2(k+1)(1-z)^{N-1}(1+z)^{M-1}P_{k+1}^{(N-1,M-1)}(z),$ (4.7)
which also makes its appearance in connection with the shape invariance of the TDPT potential[35].
Then
$\psi_{k}^{\left(-n-1\right)}\left(x;N,M;\lambda\right)=\frac{\psi_{0}(x;N-1,M-1)}{p_{n}^{(-N,-M)}(z;\lambda)}Q_{k}^{\left(n\right)}\left(z;N,M;\lambda\right),$ (4.8)
where the $Q_{k}^{\left(n\right)}$’s are given by $Q_{-n-1}^{\left(n\right)}\left(z;N,M;\lambda\right)=1$ and, for $k\geq 0$,
$Q_{k}^{\left(n\right)}\left(z;N,M;\lambda\right)$ $=$ $\left|\begin{array}[]{cc}p_{n}^{(-N,-M)}(z;\lambda)&(1-z^{2})P_{k}^{(N,M)}(z)\\ np_{n-1}^{\left(-N+1,-M+1\right)}\left(z;a_{n}^{(N,M)}\lambda\right)&-2(k+1)P_{k+1}^{(N-1,M-1)}(z)\end{array}\right|$ (4.12)
$=$ $-2(k+1)P_{k+1}^{(N-1,M-1)}(z)p_{n}^{(-N,-M)}(z;\lambda)$
$-n(1-z^{2})P_{k}^{(N,M)}(z)p_{n-1}^{\left(-N+1,-M+1\right)}\left(z;\frac{M+N-n-1}{n}\lambda\right),$
The set of $Q_{k}^{(n)}(z;N,M;\lambda)$, $k=-n-1,0,1,2,\ldots$, is made of orthogonal polynomials on ]1,1[]-1,1[] - 1 , 1 [ with respect to the measure
$\mu_{n}^{\left(-N,-M\right)}\left(z;\lambda\right)=\frac{(1-z)^{N-1}(1+z)^{M-1}}{\left(p_{n}^{\left(-N,-M\right)}\left(z;\lambda\right)\right)^{2}}.$ (4.13)
V TWO-STEP REGULAR EXTENSIONS OF THE TDPT POTENTIAL
A Regularity conditions
Consider now a two-step chain of DT based on PJ seed functions $\psi_{-n_{1}-1}\left(x;N,M;\lambda_{1}\right)$ and $\psi_{-n_{2}-1}\left(x;N,M;\lambda_{2}\right)$,
$V\left(x;N,M\right)\overset{A(\psi_{-n_{1}-1})}{\longrightarrow}V^{\left(-n_{1}-1\right)}\left(x;N,M;\lambda_{1}\right)\overset{A\left(\psi_{-n_{2}-1}^{\left(-n_{1}-1\right)}\right)}{\longrightarrow}V^{\left(-n_{1}-1,-n_{2}-1\right)}\left(x;N,M;\lambda_{1},\lambda_{2}\right),$ (5.1)
where
$\psi_{-n_{2}-1}^{\left(-n_{1}-1\right)}\left(x;N,M;\lambda_{1},\lambda_{2}\right)=\frac{W\left(\psi_{-n_{1}-1}(x;N,M;\lambda_{1}),\psi_{-n_{2}-1}(x;N,M;\lambda_{2})\mid x\right)}{\psi_{-n_{1}-1}\left(x;N,M;\lambda_{1}\right)}.$ (5.2)
We are interested in chains for which both extensions $V^{\left(-n_{1}-1\right)}$ and $V^{\left(-n_{1}-1,-n_{2}-1\right)}$ are regular. For such a purpose, we have first to choose $\max\left(N,M\right)\leq n_{1}<N+M$ and $\lambda_{1}$ in order to satisfy the regularity condition (4.1), which ensures that the seed function $\psi_{-n_{1}-1}$ has no node on $]0,\pi/2[$. We then choose $n_{2}$ in order that the seed function of the second DT $\psi_{-n_{2}-1}^{\left(-n_{1}-1\right)}$ be in the disconjugacy sector of $V^{\left(-n_{1}-1\right)}$, namely
$E_{-n_{1}-1}\left(N,M\right)-E_{-n_{2}-1}\left(N,M\right)=4(n_{1}-n_{2})\left(n_{1}+n_{2}-N-M+1\right)>0.$ (5.3)
This is achieved if
$\max\left(N,M\right)\leq n_{2}<n_{1}<N+M.$ (5.4)
$V^{\left(-n_{1}-1,-n_{2}-1\right)}\left(x;N,M;\lambda_{1},\lambda_{2}\right)$ is then regular when $\psi_{-n_{2}-1}^{\left(-n_{1}-1\right)}\left(x;N,M;\lambda_{1},\lambda_{2}\right)$ is nodeless on $]0,\pi/2[$. Since $\psi_{-n_{1}-1}$ does not change its sign on this interval, this corresponds to
$\mathrm{sign}\left(W\left(\psi_{-n_{1}-1},\psi_{-n_{2}-1}\mid 0\right)\right)=\mathrm{sign}\left(W\left(\psi_{-n_{1}-1},\psi_{-n_{2}-1}\mid\pi/2\right)\right).$ (5.5)
But we have
$W\left(\psi_{-n_{1}-1},\psi_{-n_{2}-1}\mid x\right)$ $=$ $-2(1-z)^{-N+1}(1+z)^{-M+1}$ (5.6)
×$\times W\left(p_{n_{1}}^{\left(-N,-M\right)}\left(z;\lambda_{1}\right),p_{n_{2}}^{\left(-N,-M\right)}\left(z;\lambda_{2}\right)\mid z\right),$
from which we deduce that the preceding condition can be rewritten as
$\mathrm{sign}\left(W\left(p_{n_{1}}^{\left(-N,-M\right)}\left(z;\lambda_{1}\right),p_{n_{2}}^{\left(-N,-M\right)}\left(z;\lambda_{2}\right)\mid 1\right)\right)$
$=\mathrm{sign}\left(W\left(p_{n_{1}}^{\left(-N,-M\right)}\left(z;\lambda_{1}\right),p_{n_{2}}^{\left(-N,-M\right)}\left(z;\lambda_{2}\right)\mid-1\right)\right).$ (5.7)
On using Eqs. (3.16), (3.18), (3.19), and (3.25), we obtain
$W\left(p_{n_{1}}^{\left(-N,-M\right)}\left(z;\lambda_{1}\right),p_{n_{2}}^{\left(-N,-M\right)}\left(z;\lambda_{2}\right)\mid-1\right)$ (5.10)
$=$ $\left|\begin{array}[]{cc}\left(-1\right)^{n_{1}}b_{n_{1}}^{\left(N,M\right)}\lambda_{1}&\left(-1\right)^{n_{2}}b_{n_{2}}^{\left(N,M\right)}\lambda_{2}\\ n_{1}\left(-1\right)^{n_{1}-1}b_{n_{1}-1}^{\left(N-1,M-1\right)}a_{n_{1}}^{\left(N,M\right)}\lambda_{1}&n_{2}\left(-1\right)^{n_{2}-1}b_{n_{2}-1}^{\left(N-1,M-1\right)}a_{n_{2}}^{\left(N,M\right)}\lambda_{2}\end{array}\right|$
$=$ $\left(-1\right)^{n_{1}+n_{2}}\frac{b_{n_{1}}^{\left(N,M\right)}b_{n_{2}}^{\left(N,M\right)}}{2(M-1)}(n_{2}-n_{1})\left(n_{1}+n_{2}-N-M+1\right)\lambda_{1}\lambda_{2},$ (5.11)
or, with Eq. (5.3),
$W\left(p_{n_{1}}^{\left(-N,-M\right)}\left(z;\lambda_{1}\right),p_{n_{2}}^{\left(-N,-M\right)}\left(z;\lambda_{2}\right)\mid-1\right)$ (5.12)
$=$ $\left(-1\right)^{n_{1}+n_{2}-1}\frac{b_{n_{1}}^{\left(N,M\right)}b_{n_{2}}^{\left(N,M\right)}}{8(M-1)}\left(E_{-n_{1}-1}\left(N,M\right)-E_{-n_{2}-1}\left(N,M\right)\right)\lambda_{1}\lambda_{2}.$
On the other hand (see Eqs. (3.16), (3.18), (3.20), and (3.25)), $W\left(p_{n_{1}}^{\left(-N,-M\right)}\left(z;\lambda_{1}\right),p_{n_{2}}^{\left(-N,-M\right)}\left(z;\lambda_{2}\right)\mid+1\right)$ can be written as
$W\left(p_{n_{1}}^{\left(-N,-M\right)}\left(z;\lambda_{1}\right),p_{n_{2}}^{\left(-N,-M\right)}\left(z;\lambda_{2}\right)\mid+1\right)$ (5.15)
$=$ $\left|\begin{array}[]{cc}p_{n_{1}}^{\left(-N,-M\right)}\left(1;\lambda_{1}\right)&p_{n_{2}}^{\left(-N,-M\right)}\left(1;\lambda_{2}\right)\\ n_{1}p_{n_{1}-1}^{\left(-N+1,-M+1\right)}\left(1;a_{n_{1}}^{\left(N,M\right)}\lambda_{1}\right)&n_{2}p_{n_{2}-1}^{\left(-N+1,-M+1\right)}\left(1;a_{n_{2}}^{\left(N,M\right)}\lambda_{2}\right)\end{array}\right|$
$=$ $\left(-1\right)^{n_{1}+n_{2}}\frac{b_{n_{1}}^{\left(N,M\right)}b_{n_{2}}^{\left(N,M\right)}}{2(N-1)}\times A,$ (5.16)
where (see Eqs. (3.19), (3.21), and (3.24))
$A$ $=$ $\left|\begin{array}[]{cc}\left(-1\right)^{n_{1}-N+1}\lambda_{1}+\lambda_{n_{1}}^{\left(N,M\right)}&\left(-1\right)^{n_{2}-N+1}\lambda_{2}+\lambda_{n_{2}}^{\left(N,M\right)}\\ n_{1}^{2}a_{n_{1}}^{\left(N,M\right)}\left(\left(-1\right)^{n_{1}-N+1}\lambda_{1}+\lambda_{n_{1}}^{\left(N,M\right)}\right)&n_{2}^{2}a_{n_{2}}^{\left(N,M\right)}\left(\left(-1\right)^{n_{2}-N+1}\lambda_{2}+\lambda_{n_{2}}^{\left(N,M\right)}\right)\end{array}\right|$ (5.20)
$=$ $\frac{\left(\left(-1\right)^{n_{1}-N+1}\lambda_{1}+\lambda_{n_{1}}^{\left(N,M\right)}\right)\left(\left(-1\right)^{n_{2}-N+1}\lambda_{2}+\lambda_{n_{2}}^{\left(N,M\right)}\right)}{4}$
×$\times\left(E_{-n_{1}-1}\left(N,M\right)-E_{-n_{2}-1}\left(N,M\right)\right).$
Consequently
$W\left(p_{n_{1}}^{\left(-N,-M\right)}\left(z;\lambda_{1}\right),p_{n_{2}}^{\left(-N,-M\right)}\left(z;\lambda_{2}\right)\mid+1\right)$ (5.21)
$=$ $\frac{b_{n_{1}}^{\left(N,M\right)}b_{n_{2}}^{\left(N,M\right)}}{8(N-1)}\left(\lambda_{1}+\left(-1\right)^{n_{1}-N+1}\lambda_{n_{1}}^{\left(N,M\right)}\right)\left(\lambda_{2}+\left(-1\right)^{n_{2}-N+1}\lambda_{n_{2}}^{\left(N,M\right)}\right)$
×$\times\left(E_{-n_{1}-1}\left(N,M\right)-E_{-n_{2}-1}\left(N,M\right)\right).$
By comparing Eq. (5.12) and Eq. (5.21), the regularity condition (5.7) for the two-step extension $V^{\left(-n_{1}-1,-n_{2}-1)\right)}\left(x;N,M;\lambda_{1},\lambda_{2}\right)$ can be written as
$\mathrm{sign}\left(\lambda_{1}\lambda_{2}\right)=\mathrm{sign}\left(\left(-1\right)^{n_{1}+n_{2}-1}\left(\lambda_{1}+\left(-1\right)^{n_{1}-N+1}\lambda_{n_{1}}^{\left(N,M\right)}\right)\left(\lambda_{2}+\left(-1\right)^{n_{2}-N+1}\lambda_{n_{2}}^{\left(N,M\right)}\right)\right).$ (5.22)
Since the regularity condition for the first extension $V^{-n_{1}-1}\left(x;N,M;\lambda_{1}\right)$ (see [24]) implies that
$\mathrm{sign}\left(\lambda_{1}\right)=\mathrm{sign}\left(\left(-1\right)^{n_{1}}g_{n_{1}}^{\left(N,M\right)}\left(\lambda_{1}\right)\right)=\mathrm{sign}\left(\left(-1\right)^{n_{1}-N-M+1}\left(\lambda_{1}+\left(-1\right)^{n_{1}-N+1}\lambda_{n_{1}}^{\left(N,M\right)}\right)\right),$ (5.23)
Eq. (5.22) simply becomes
$\mathrm{sign}\left(\lambda_{2}\right)=\mathrm{sign}\left(\left(-1\right)^{n_{2}-N-M}\lambda_{2}+\left(-1\right)^{M-1}\lambda_{n_{2}}^{\left(N,M\right)}\right).$ (5.24)
Consider first the case where $M$, $n_{1}-N$, and $n_{2}-N$ are all even. The regularity of $V^{\left(-n_{1}-1\right)}\left(x;N,M;\lambda_{1}\right)$ necessitates $0<\lambda_{1}<\lambda_{n_{1}}^{\left(N,M\right)}$. Then condition (5.24) becomes
$\mathrm{sign}\left(\lambda_{2}\right)=\mathrm{sign}\left(\lambda_{2}-\lambda_{n_{2}}^{\left(N,M\right)}\right).$ (5.25)
This condition is always achieved when $\lambda_{2}<0$ and if $\lambda_{2}>0$ provided
$\lambda_{2}>\lambda_{n_{2}}^{\left(N,M\right)}.$ (5.26)
In the same manner, we find the other regularity conditions. Consequently, we have the regularity conditions mentioned below:
$\left\{\begin{array}[]{c}(M,n_{1}-N,n_{2}-N)\in\left(2\mathbb{N}\right)^{3}:0<\lambda_{1}<\lambda_{n_{1}}^{\left(N,M\right)},\lambda_{2}<0\text{ or }\lambda_{2}>\lambda_{n_{2}}^{(N,M)};\\[5.69046pt] (M,n_{1}-N,n_{2}-N)\in\left(2\mathbb{N}\right)^{2}\times(2\mathbb{N}+1):0<\lambda_{1}<\lambda_{n_{1}}^{\left(N,M\right)},-\lambda_{n_{2}}^{(N,M)}<\lambda_{2}<0;\\[5.69046pt] (M,n_{1}-N,n_{2}-N)\in 2\mathbb{N}\times(2\mathbb{N}+1)\times 2\mathbb{N}:\lambda_{1}<-\lambda_{n_{1}}^{(N,M)}\text{ or }\lambda_{1}>0,\lambda_{2}<0\\ \text{ or }\lambda_{2}>\lambda_{n_{2}}^{(N,M)};\\[5.69046pt] (M,n_{1}-N,n_{2}-N)\in 2\mathbb{N}\times(2\mathbb{N}+1)^{2}:\lambda_{1}<-\lambda_{n_{1}}^{(N,M)}\text{ or }\lambda_{1}>0,\\ -\lambda_{n_{2}}^{(N,M)}<\lambda_{2}<0;\\[5.69046pt] (M,n_{1}-N,n_{2}-N)\in(2\mathbb{N}+1)\times(2\mathbb{N})^{2}:\lambda_{1}<0\text{ or }\lambda_{1}>\lambda_{n_{1}}^{(N,M)},\\ 0<\lambda_{2}<\lambda_{n_{2}}^{(N,M)};\\[5.69046pt] (M,n_{1}-N,n_{2}-N)\in(2\mathbb{N}+1)\times 2\mathbb{N}\times(2\mathbb{N}+1):\lambda_{1}<0\text{ or }\lambda_{1}>\lambda_{n_{1}}^{(N,M)},\\ \lambda_{2}<-\lambda_{n_{2}}^{(N,M)}\text{ or }\lambda_{2}>0;\\[5.69046pt] (M,n_{1}-N,n_{2}-N)\in(2\mathbb{N}+1)^{2}\times 2\mathbb{N}:-\lambda_{n_{1}}^{(N,M)}<\lambda_{1}<0,0<\lambda_{2}<\lambda_{n_{2}}^{(N,M)};\\[5.69046pt] (M,n_{1}-N,n_{2}-N)\in(2\mathbb{N}+1)^{3}:-\lambda_{n_{1}}^{(N,M)}<\lambda_{1}<0,\lambda_{2}<-\lambda_{n_{2}}^{(N,M)}\text{ or }\lambda_{2}>0.\end{array}\right.$ (5.27)
B Rationally-extended potential
If $\lambda_{1}$ and $\lambda_{2}$ satisfy conditions (5.27), then $\psi_{-n_{1}-1}\left(x;N,M;\lambda_{1}\right)$ and $\psi_{-n_{2}-1}\left(x;N,M;\lambda_{2}\right)$ can be used as seed functions to build a two-step state-adding chain of DT, which generates a regular rational extension of $V\left(x;N-2,M-2\right)$,
$V^{-n_{1}-1,-n_{2}-1}(x;N,M;\lambda_{1},\lambda_{2})$
$\quad=V(x;N,M)-2\frac{d^{2}}{dx^{2}}\log W\left(\psi_{-n_{1}-1}(x;N,M;\lambda_{1}),\psi_{-n_{2}-1}(x;N,M;\lambda_{2})\mid x\right)$
$\quad=V(x;N,M)-2\frac{d^{2}}{dx^{2}}\log\left(\psi_{-1}^{2}(x;N,M)\frac{dz}{dx}\right)$
$\qquad-2\frac{d^{2}}{dx^{2}}\log W\left(p_{n_{1}}^{(-N,-M)}(z;\lambda_{1}),p_{n_{2}}^{(-N,-M)}(z;\lambda_{2})\mid z\right)$
$\quad=V(x;N-2,M-2)+E_{-2}(N,M)$
$\qquad-2\frac{d^{2}}{dx^{2}}\log W\left(p_{n_{1}}^{(-N,-M)}(z;\lambda_{1}),p_{n_{2}}^{(-N,-M)}(z;\lambda_{2})\mid z\right),$ (5.28)
where use has been made of (2.5) as well as of the relation
$\psi_{-1}^{2}(z;N,M)\frac{dz}{dx}\propto\psi_{-1}(z;N,M)\psi_{-1}(z;N-1,M-1),$ (5.29)
the shape invariance property of the TDPT potential,
$V(x;N,M)-2\frac{d^{2}}{dx^{2}}\log\psi_{-1}(z;N,M)=V(x;N-1,M-1)+E_{-1}(N,M),$ (5.30)
and the relation
$E_{-2}(N,M)=E_{-1}(N,M)+E_{-1}(N-1,M-1).$ (5.31)
It only remains to make the change of variable from $x$ to $z$ to get the final result
$V^{(-n_{1}-1,-n_{2}-1)}(x;N,M;\lambda_{1},\lambda_{2})$
$\quad=V(x;N-2,M-2)+E_{-2}(N,M)$
$\qquad-8(1-z^{2})\frac{d^{2}}{dz^{2}}\log W\left(p_{n_{1}}^{(-N,-M)}(z;\lambda_{1}),p_{n_{2}}^{(-N,-M)}(z;\lambda_{2})\mid z\right)$
$\qquad+8z\frac{d}{dz}\log W\left(p_{n_{1}}^{(-N,-M)}(z;\lambda_{1}),p_{n_{2}}^{(-N,-M)}(z;\lambda_{2})\mid z\right).$ (5.32)
For $V^{(-n_{1}-1,-n_{2}-1)}(x;N,M;\lambda_{1},\lambda_{2})$ to be a confining potential on $]0,\pi/2[$, we need to impose this property to $V(x;N-2,M-2)$. This is achieved for $N,M\geq 3$. For such a choice and provided $n_{1}$ and $n_{2}$ satisfy Eq. (5.4), $V^{(-n_{1}-1,-n_{2}-1)}(x;N,M;\lambda_{1},\lambda_{2})$ is strongly repulsive in both 0 and $\pi/2$, so that in each extremity only one basis solution is quadratically integrable and the corresponding Hamiltonian is essentially self-adjoint [33].
The corresponding eigenstates are given by
$\psi_{k}^{(-n_{1}-1,-n_{2}-1)}(x;N,M;\lambda_{1},\lambda_{2})$
$\quad=\frac{W(\psi_{-n_{1}-1}(x;N,M;\lambda_{1},\lambda_{2}),\psi_{-n_{2}-1}(x;N,M;\lambda_{1},\lambda_{2}),\psi_{k}(x;N,M)\mid x)}{W(\psi_{-n_{1}-1}(x;N,M;\lambda_{1},\lambda_{2}),\psi_{-n_{2}-1}(x;N,M;\lambda_{1},\lambda_{2})\mid x)},$ (5.33)
for $k=0,1,2,\ldots$, as well as
$\psi_{-n_{1}-1}^{(-n_{1}-1,-n_{2}-1)}(x;N,M;\lambda_{1},\lambda_{2})$
$\quad=\frac{\psi_{-n_{2}-1}(x;N,M;\lambda_{1},\lambda_{2})}{W(\psi_{-n_{1}-1}(x;N,M;\lambda_{1},\lambda_{2}),\psi_{-n_{2}-1}(x;N,M;\lambda_{1},\lambda_{2})\mid x)},$ (5.34)
and
$\psi_{-n_{2}-1}^{(-n_{1}-1,-n_{2}-1)}(x;N,M;\lambda_{1},\lambda_{2})$
$\quad=\frac{\psi_{-n_{1}-1}(x;N,M;\lambda_{1},\lambda_{2})}{W(\psi_{-n_{1}-1}(x;N,M;\lambda_{1},\lambda_{2}),\psi_{-n_{2}-1}(x;N,M;\lambda_{1},\lambda_{2})\mid x)}.$ (5.35)
The corresponding energies are $E_{k}(N,M)$ ($k=0,1,2,\ldots$), $E_{-n_{1}-1}(N,M)$, and $E_{-n_{2}-1}(N,M)$, respectively.
On using (2.5) again, as well as derivative properties of Jacobi and para-Jacobi polynomials, Eqs. (5.33), (5.34), and (5.35) can be rewritten as
$\psi_{k}^{(-n_{1}-1,-n_{2}-1)}(x;N,M;\lambda_{1},\lambda_{2})\propto\psi_{0}(x;N-2,M-2)\frac{Q_{k}^{(n_{1},n_{2})}(z;N,M;\lambda_{1},\lambda_{2})}{T_{n_{1}n_{2}}^{(-N,-M)}(z;\lambda_{1},\lambda_{2})},$ (5.36)
where
$Q_{k}^{(n_{1},n_{2})}(z;N,M;\lambda_{1},\lambda_{2})=4(k+1)(k+2)T_{n_{1}n_{2}}^{(-N,-M)}(z;\lambda_{1},\lambda_{2})P_{k+2}^{(N-2,M-2)}(z)$
$\quad-2(k+1)U_{n_{1}n_{2}}^{(-N,-M)}(z;\lambda_{1},\lambda_{2})(1-z^{2})P_{k+1}^{(N-1,M-1)}(z)$
$\quad+V_{n_{1}n_{2}}^{(-N,-M)}(z;\lambda_{1},\lambda_{2})(1-z^{2})^{2}P_{k}^{(N,M)}(z),\qquad k=0,1,2,\ldots,$ (5.37)
$Q_{-n_{1}-1}^{(n_{1},n_{2})}(z;N,M;\lambda_{1},\lambda_{2})=p_{n_{2}}^{(-N,-M)}(z;\lambda_{2}),$ (5.38)
$Q_{-n_{2}-1}^{(n_{1},n_{2})}(z;N,M;\lambda_{1},\lambda_{2})=p_{n_{1}}^{(-N,-M)}(z;\lambda_{1}).$ (5.39)
In these equations, we have defined
$T_{n_{1}n_{2}}^{(-N,-M)}(z;\lambda_{1},\lambda_{2})$
$\quad=n_{2}p_{n_{1}}^{(-N,-M)}(z;\lambda_{1})p_{n_{2}-1}^{(-N+1,-M+1)}\left(z;a_{n_{2}}^{(N,M)}\lambda_{2}\right)$
$\qquad-n_{1}p_{n_{1}-1}^{(-N+1,-M+1)}\left(z;a_{n_{1}}^{(N,M)}\lambda_{1}\right)p_{n_{2}}^{(-N,-M)}(z;\lambda_{2}),$ (5.40)
$U_{n_{1}n_{2}}^{(-N,-M)}(z;\lambda_{1},\lambda_{2})$
$\quad=n_{2}(n_{2}-1)p_{n_{1}}^{(-N,-M)}(z;\lambda_{1})p_{n_{2}-2}^{(-N+2,-M+2)}\left(z;a_{n_{2}-1}^{(N-1,M-1)}a_{n_{2}}^{(N,M)}\lambda_{2}\right)$
$\qquad-n_{1}(n_{1}-1)p_{n_{1}-2}^{(-N+2,-M+2)}\left(z;a_{n_{1}-1}^{(N-1,M-1)}a_{n_{1}}^{(N,M)}\lambda_{1}\right)p_{n_{2}}^{(-N,-M)}(z;\lambda_{2}),$ (5.41)
$V_{n_{1}n_{2}}^{(-N,-M)}(z;\lambda_{1},\lambda_{2})$
$\quad=n_{1}n_{2}(n_{2}-1)p_{n_{1}-1}^{(-N+1,-M+1)}\left(z;a_{n_{1}}^{(N,M)}\lambda_{1}\right)p_{n_{2}-2}^{(-N+2,-M+2)}\left(z;a_{n_{2}-1}^{(N-1,M-1)}a_{n_{2}}^{(N,M)}\lambda_{2}\right)$
$\qquad-n_{1}(n_{1}-1)n_{2}p_{n_{1}-2}^{(-N+2,-M+2)}\left(z;a_{n_{1}-1}^{(N-1,M-1)}a_{n_{1}}^{(N,M)}\lambda_{1}\right)$
×$\qquad\times p_{n_{2}-1}^{(-N+1,-M+1)}\left(z;a_{n_{2}}^{(N,M)}\lambda_{2}\right).$ (5.42)
Due to the orthogonality properties of the $\psi_{k}^{(-n_{1}-1,-n_{2}-1)}(x;N,M;\lambda_{1},\lambda_{2})$, we deduce that the $Q_{k}^{(n_{1},n_{2})}(z;N,M;\lambda_{1},\lambda_{2})$, $k=-n_{1}-1,-n_{2}-1,0,1,2,\ldots$, are orthogonal polynomials on ]1,1[]-1,1[] - 1 , 1 [ with respect to the measure
$\mu_{n_{1}n_{2}}^{(-N,-M)}(z;\lambda_{1},\lambda_{2})=\frac{(1-z)^{N-2}(1+z)^{M-2}}{\left(T_{n_{1}n_{2}}^{(-N,-M)}(z;\lambda_{1},\lambda_{2})\right)^{2}}.$ (5.43)
C Explicit examples
Consider for instance the case where $N=M=3$. The seed function indices $\left(n_{1},n_{2}\right)$ of the two-step chain can be chosen in the set ($63\leq n_{2}<n_{1}<$ 6), so that $(n_{1},n_{2})\in\left\{\left(4,3\right),(5,4),(5,3)\right\}$. As for the possible values of the regularity parameters $\lambda_{n_{i}}^{\left(N,M\right)},\ i=1,2$, we get from Eq. (3.22)
$\lambda_{3}^{\left(3,3\right)}=3,\ \lambda_{4}^{\left(3,3\right)}=2,\ \lambda_{5}^{\left(3,3\right)}=1/6.$ (5.44)
From Eq. (3.13), the corresponding para-Jacobi polynomials can be written as
$\left\{\begin{array}[]{c}p_{3}^{\left(-3,-3\right)}\left(z;\lambda\right)=z^{3}+(3-2\lambda)z^{2}+3z+1-\frac{2}{3}\lambda,\\[2.84544pt] p_{4}^{\left(-3,-3\right)}\left(z;\lambda\right)=z^{4}-6z^{2}-8(\lambda+1)z-3,\\[2.84544pt] p_{5}^{\left(-3,-3\right)}\left(z;\lambda\right)=z^{5}-\frac{10}{3}z^{3}+5z+8(\frac{1}{3}-4\lambda).\end{array}\right.$ (5.45)
Then, from Eq. (5.32),
$V^{\left(-n_{1}-1,-n_{2}-1\right)}\left(x;3,3;\lambda_{1},\lambda_{2}\right)$ $=$ $V(x;1,1)+E_{-2}(3,3)$ (5.46)
${}+\Delta V^{\left(-n_{1}-1,-n_{2}-1\right)}\left(x;3,3;\lambda_{1},\lambda_{2}\right),$
with
$V(x;1,1)=\frac{3}{1-z^{2}}-9,\qquad E_{-2}(3,3)=-40,$ (5.47)
and
$\Delta V^{\left(-n_{1}-1,-n_{2}-1\right)}\left(x;3,3;\lambda_{1},\lambda_{2}\right)=8z\frac{d}{dz}\log W\left(p_{n_{1}}^{\left(-3,-3\right)}\left(z;\lambda_{1}\right),p_{n_{2}}^{\left(-3,-3\right)}\left(z;\lambda_{2}\right)\mid z\right)$
$\qquad{}-8(1-z^{2})\frac{d^{2}}{dz^{2}}\log W\left(p_{n_{1}}^{\left(-3,-3\right)}\left(z;\lambda_{1}\right),p_{n_{2}}^{\left(-3,-3\right)}\left(z;\lambda_{2}\right)\mid z\right).$ (5.48)
If $n_{1}=4$ and $n_{2}=3$, for instance,
$\Omega\left(z;\lambda_{1},\lambda_{2}\right)$
$\qquad=W\left(p_{n_{1}}^{\left(-3,-3\right)}\left(z;\lambda_{1}\right),p_{n_{2}}^{\left(-3,-3\right)}\left(z;\lambda_{2}\right)\mid z\right)$
$\qquad=\left|\begin{array}[]{cc}z^{4}-6z^{2}-8(\lambda_{1}+1)z-3&z^{3}+(3-2\lambda_{2})z^{2}+3z+1-\frac{2}{3}\lambda_{2}\\ 4z^{3}-12z-8(\lambda_{1}+1)&3z^{2}+2(3-2\lambda_{2})z+3\end{array}\right|$ (5.51)
$\qquad=-z^{6}+\left(4\lambda_{2}-6\right)z^{5}-15z^{4}+\left(\frac{8}{3}\lambda_{2}-16\lambda_{1}-20\right)z^{3}$
$\qquad\quad+(16\lambda_{2}\lambda_{1}-24\lambda_{1}+16\lambda_{2}-15)z^{2}+\left(4\lambda_{2}-6\right)z+8\lambda_{1}-\frac{16}{3}\lambda_{2}$
$\qquad\quad-\frac{16}{3}\lambda_{1}\lambda_{2}-1$ (5.52)
and
$V^{\left(-5,-4\right)}\left(z;3,3;\lambda_{1},\lambda_{2}\right)$
$\qquad=\frac{3}{1-z^{2}}-17+32(1-z^{2})\left(\frac{A\left(z;\lambda_{1},\lambda_{2}\right)}{\Omega\left(z;\lambda_{1},\lambda_{2}\right)}\right)^{2}-16\frac{B\left(z;\lambda_{1},\lambda_{2}\right)}{\Omega\left(z;\lambda_{1},\lambda_{2}\right)},$ (5.53)
with
$\left\{\begin{array}[]{c}A\left(z;\lambda_{1},\lambda_{2}\right)=3z^{5}+\left(15-10\lambda_{2}\right)z^{4}+30z^{3}+\left(24\lambda_{1}-4\lambda_{2}+30\right)z^{2}\\ +(24\lambda_{1}-16\lambda_{2}-16\lambda_{2}\lambda_{1}+15)z-\left(2\lambda_{2}-3\right),\\[2.84544pt] B\left(z;\lambda_{1},\lambda_{2}\right)=18z^{6}-\left(50\lambda_{2}-75\right)z^{5}+105z^{4}+\left(28\lambda_{2}+72\lambda_{1}+30\right)z^{3}\\ +(48\lambda_{1}-32\lambda_{2}-32\lambda_{2}\lambda_{1}-60)z^{2}+\left(6\lambda_{2}-48\lambda_{1}-57\right)z\\ +(16\lambda_{2}-24\lambda_{1}+16\lambda_{1}\lambda_{2}-15).\end{array}\right.$ (5.54)
As for the regularity conditions, they are given by
$\left\{\begin{array}[]{c}-2<\lambda_{1}<0,\\ 0<\lambda_{2}<3.\end{array}\right.$ (5.55)
In Fig. 1, potential (5.53) is plotted in terms of $z$ for $\lambda_{1}=-1$ and $\lambda_{2}=1$. It is also displayed in terms of $z$ and $t=\lambda_{1}$ for $\lambda_{2}=1$ or in terms of $z$ and $\mu=\lambda_{2}$ for $\lambda_{1}=-1$ in Figs. 2 and 3, respectively.
Figure 1: Potential $V^{(-5,-4)}(z;3,3;-1,1)$ in terms of $z$ for $-1<z<1$.
Figure 2: Potential $V^{(-5,-4)}(z;3,3;t,1)$ in terms of $z$ and $t$ for $-1<z<1$ and $-2<t<0$.
Figure 3: Potential $V^{(-5,-4)}(z;3,3;-1,\mu)$ in terms of $z$ and $\mu$ for $-1<z<1$ and $30<\mu<$ 3.
VI $m$-STEP REGULAR EXTENSIONS OF THE TDPT POTENTIAL
A Regularity conditions
Consider now an $m$-step chain $N_{m}=(-n_{1}-1,-n_{2}-1,\ldots,-n_{m}-1)$ of DT based on PJ seed functions $\psi_{-n_{1}-1}(x;N,M;\lambda_{1})$, $\psi_{-n_{2}-1}(x;N,M;\lambda_{2})$, …, $\psi_{-n_{m}-1}(x;N,M;\lambda_{m})$,
$V\left(x;N,M\right)\overset{A(\psi_{-n_{1}-1})}{\longrightarrow}V^{\left(-n_{1}-1\right)}\left(x;N,M;\lambda_{1}\right)\overset{A\left(\psi^{(-n_{1}-1)}_{-n_{2}-1}\right)}{\longrightarrow}\cdots$
$\overset{A\left(\psi^{(N_{m}-2)}_{-n_{m-1}-1}\right)}{\longrightarrow}V^{(N_{m-1})}(x;N,M;\lambda_{1},\ldots,\lambda_{m-1})\overset{A\left(\psi^{(N_{m-1})}_{-n_{m}-1}\right)}{\longrightarrow}V^{(N_{m})}(x;N,M;\lambda_{1},\ldots,\lambda_{m}),$ (6.1)
where
$\psi^{(N_{m-1})}_{-n_{m}-1}(x;N,M;\lambda_{1},\ldots,\lambda_{m})$
$\quad=\frac{W(\psi_{-n_{1}-1}(x;N,M;\lambda_{1}),\psi_{-n_{2}-1}(x;N,M;\lambda_{2}),\ldots,\psi_{-n_{m}-1}(x;N,M;\lambda_{m})\mid x)}{W(\psi_{-n_{1}-1}(x;N,M;\lambda_{1}),\psi_{-n_{2}-1}(x;N,M;\lambda_{2}),\ldots,\psi_{-n_{m-1}-1}(x;N,M;\lambda_{m-1})\mid x)}.$ (6.2)
On assuming that the extension obtained at the $(m-1)$th step is regular, which imposes that $\max\left(N,M\right)\leq n_{m-1}<\cdots<n_{2}<n_{1}<N+M$ and definite conditions on $\lambda_{1},\lambda_{2},\ldots,\lambda_{m-1}$, we want to impose that the same is true at the $m$th step. For this to occur, $\psi^{(N_{m-1})}_{-n_{m}-1}$ must be in the disconjugacy sector of $V^{(N_{m-1})}$, i.e., $\max\left(N,M\right)\leq n_{m}<n_{m-1}<\cdots<n_{1}<N+M$, and, in addition, it must not change sign on $]0,\pi/2[$, which leads to the condition
$\mathrm{sign}(W(\psi_{-n_{1}-1},\psi_{-n_{2}-1},\ldots,\psi_{-n_{m}-1}\mid 0))$
$\quad=\mathrm{sign}(W(\psi_{-n_{1}-1},\psi_{-n_{2}-1},\ldots,\psi_{-n_{m}-1}\mid\pi/2)).$ (6.3)
Since standard properties of Wronskians (2.5) lead to
$W(\psi_{-n_{1}-1},\psi_{-n_{2}-1},\ldots,\psi_{-n_{m}-1}\mid x)\propto(1-z)^{\frac{m}{2}\left(-N+\frac{m}{2}\right)}(1+z)^{\frac{m}{2}\left(-M+\frac{m}{2}\right)}$
×$\qquad\times W\left(p_{n_{1}}^{(-N,-M)}(z;\lambda_{1}),p_{n_{2}}^{(-N,-M)}(z;\lambda_{2}),\ldots,p_{n_{m}}^{(-N,-M)}(z;\lambda_{m})\mid z\right),$ (6.4)
condition (6.3) can be rewritten as
${\rm sign}\left(W\left(p_{n_{1}}^{\left(-N,-M\right)}\left(z;\lambda_{1}\right),\ldots,p_{n_{m}}^{\left(-N,-M\right)}\left(z;\lambda_{m}\right)\mid 1\right)\right)$
$\quad={\rm sign}\left(W\left(p_{n_{1}}^{\left(-N,-M\right)}\left(z;\lambda_{1}\right),\ldots,p_{n_{m}}^{\left(-N,-M\right)}\left(z;\lambda_{m}\right)\mid-1\right)\right).$ (6.5)
The derivation rule for the para-Jacobi polynomials (3.18) leads to
$\frac{d^{k}}{dz^{k}}p_{n_{j}}^{\left(-N,-M\right)}\left(z;\lambda\right)=\left(n_{j}\right)_{\underline{k}}p_{n_{j}-k}^{\left(-N+k,-M+k\right)}\left(z;A_{j}^{(k)}\lambda\right),$ (6.6)
where (cf Eq. (3.19))
$A_{j}^{(k)}=\prod_{i=0}^{k-1}a_{n_{j}-i}^{\left(N-i,M-i\right)}=\frac{\left(N+M-n_{i}-1\right)_{\underline{k}}}{\left(n\right)_{\underline{k}}},$ (6.7)
$(x)_{\underline{p}}$ being the falling factorial
$(x)_{\underline{l}}=x(x-1)...(x-l+1).$ (6.8)
This gives
$W\left(p_{n_{1}}^{(-N,-M)}(z;\lambda_{1}),p_{n_{2}}^{(-N,-M)}(z;\lambda_{2}),\ldots,p_{n_{m}}^{(-N,-M)}(z;\lambda_{m})\mid z\right)=\det\left[\mathbf{R}\left(z;\lambda_{1},...,\lambda_{m}\right)\right],$ (6.9)
where the elements of the matrix are given by
$\mathbf{R}_{i,j}\left(z;\lambda_{1},...,\lambda_{m}\right)=\left(n_{j}\right)_{\underline{i-1}}p_{n_{j}-i+1}^{\left(-N+i-1,-M+i-1\right)}\left(z;A_{j}^{(i-1)}\lambda_{j}\right).$ (6.10)
Condition (6.3) can then be rewritten as
$\mathrm{sign}\left(\det\left[\mathbf{R}\left(1;\lambda_{1},...,\lambda_{m}\right)\right]\right)=\mathrm{sign}\left(\det\left[\mathbf{R}\left(-1;\lambda_{1},...,\lambda_{m}\right)\right]\right).$ (6.11)
Consider first the right-hand side of Eq. (6.11). By using Eqs. (3.16), (3.17), and since, from (3.25),
$b_{n-k}^{\left(N-k,M-k\right)}=\frac{\left(n\right)_{\underline{k}}}{2^{k}(M-1)_{\underline{k}}}b_{n}^{\left(N,M\right)},$ (6.12)
we obtain for $\det\left[\mathbf{R}_{i,j}\left(-1;\lambda_{1},...,\lambda_{m}\right)\right]$,
$\left|\begin{array}[]{ccc}\lambda_{1}\left(-1\right)^{n_{1}}b_{n_{1}}^{\left(N,M\right)}&...&\lambda_{m}\left(-1\right)^{n_{m}}b_{n_{m}}^{\left(N,M\right)}\\ ...&&...\\ \lambda_{1}\left(-1\right)^{n_{1}-k}A_{1}^{(k)}\frac{\left(\left(n_{1}\right)_{\underline{k}}\right)^{2}}{2^{k}(M-1)_{\underline{k}}}b_{n_{1}}^{\left(N,M\right)}&...&\lambda_{m}\left(-1\right)^{n_{m}-k}A_{m}^{(k)}\frac{\left(\left(n_{m}\right)_{\underline{k}}\right)^{2}}{2^{k}(M-1)_{\underline{k}}}b_{n_{m}}^{\left(N,M\right)}\\ ...&&...\\ \lambda_{1}\left(-1\right)^{n_{1}-m+1}A_{1}^{(m-1)}\frac{\left(\left(n_{1}\right)_{\underline{m-1}}\right)^{2}}{2^{m-1}(M-1)_{\underline{m-1}}}b_{n_{1}}^{\left(N,M\right)}&...&\lambda_{m}\left(-1\right)^{n_{m}-m+1}A_{m}^{(m-1)}\frac{\left(\left(n_{m}\right)_{\underline{m-1}}\right)^{2}}{2^{m-1}(M-1)_{\underline{m-1}}}b_{n_{m}}^{\left(N,M\right)}\end{array}\right|$ (6.18)
$=$ $\lambda_{1}...\lambda_{m}\frac{\left(-1\right)^{n_{1}+...+n_{m}+1+...+\left(m-1\right)}b_{n_{1}}^{\left(N,M\right)}...b_{n_{m}}^{\left(N,M\right)}}{2^{1+...+\left(m-1\right)}\left(M-1\right)^{m-1}\left(M-2\right)^{m-2}...(M-m+1)}D(n_{1},...,n_{m}),$ (6.19)
where
$D(n_{1},...,n_{m})$
$\quad=\left|\begin{array}[]{ccc}1&...&1\\ ...&&...\\ A_{1}^{(k)}\left(\left(n_{1}\right)_{\underline{k}}\right)^{2}&...&A_{m}^{(k)}(\left(n_{m}\right)_{\underline{k}})^{2}\\ ...&&...\\ A_{1}^{(m-1)}(\left(n_{1}\right)_{\underline{m-1}})^{2}&...&A_{m}^{(m-1)}(\left(n_{m}\right)_{\underline{m-1}})^{2}\end{array}\right|.$
Then
$\det\left[\mathbf{R}_{ij}\left(-1;\lambda_{1},...,\lambda_{m}\right)\right]$
$\quad=\lambda_{1}...\lambda_{m}\frac{\left(-1\right)^{n_{1}+...+n_{m}+m\left(m-1\right)/2}b_{n_{1}}^{\left(N,M\right)}...b_{n_{m}}^{\left(N,M\right)}}{2^{m\left(m-1\right)/2}\left(M-1\right)^{m-1}\left(M-2\right)^{m-2}...(M-m+1)}D(n_{1},...,n_{m}),$ (6.21)
with
$D(n_{1},...,n_{m})$
$\quad=\left|\begin{array}[]{ccc}1&...&1\\ ...&&...\\ \frac{n_{1}!M+N-n_{1}-1)!}{\left(n_{1}-k\right)!(M+N-n_{1}-k-1)!}&...&\frac{n_{m}!M+N-n_{m}-1)!}{\left(n_{m}-k\right)!(M+N-n_{m}-k-1)!}\\ ...&&...\\ \frac{n_{1}!M+N-n_{1}-1)!}{\left(n_{1}-m+1\right)!(M+N-n_{1}-m)!}&...&\frac{n_{m}!M+N-n_{m}-1)!}{\left(n_{m}-m+1\right)!(M+N-n_{m}-m)!}\end{array}\right|.$ (6.27)
Let us now consider $\det\left[\mathbf{R}_{ij}\left(1;\lambda_{1},...,\lambda_{m}\right)\right]$. From Eqs. (3.21) and (3.23), we have
$p_{n-k}^{\left(-N+k,-M+k\right)}\left(1;\lambda\right)=b_{n-k}^{\left(N-k,M-k\right)}\left(\left(-1\right)^{N+M-1}\lambda+\left(-1\right)^{n-M}\lambda_{n-k}^{\left(N-k,M-k\right)}\right)$ (6.28)
and from Eqs. (3.18) and (6.7), we obtain
$\frac{d^{k}}{dz^{k}}p_{n_{i}}^{\left(-N,-M\right)}\left(1;\lambda\right)=\left(n_{i}\right)_{\underline{k}}b_{n_{i}-k}^{\left(N-k,M-k\right)}\left(-1\right)^{n_{i}-M}\left(\left(-1\right)^{n_{i}-N+1}A_{i}^{(k)}\lambda+\lambda_{n_{i}-k}^{\left(N-k,M-k\right)}\right).$ (6.29)
With (cf Eq. (3.24))
$\lambda_{n_{i}-k}^{\left(N-k,M-k\right)}=A_{i}^{(k)}\lambda_{n_{i}}^{\left(N,M\right)},$ (6.30)
we therefore arrive at
$\frac{d^{k}}{dz^{k}}p_{n_{i}}^{\left(-N,-M\right)}\left(1;\lambda\right)=\left(n_{i}\right)_{\underline{k}}b_{n_{i}-k}^{\left(N-k,M-k\right)}A_{i}^{(k)}\left(-1\right)^{n_{i}-M}\left(\left(-1\right)^{n_{i}-N+1}\lambda+\lambda_{n_{i}}^{\left(N,M\right)}\right).$ (6.31)
It results that
$\det\left[\mathbf{R}\left(1;\lambda_{1},...,\lambda_{m}\right)\right].$
$\quad=\det\left(\left(n_{i}\right)_{\underline{k-1}}b_{n_{j}-k+1}^{\left(N-k+1,M-k+1\right)}A_{i}^{(k-1)}\times\left(-1\right)^{M-n_{j}}\left(\left(-1\right)^{n_{j}-N+1}\lambda_{j}+\lambda_{n_{j}}^{\left(N,M\right)}\right)\right),$ (6.32)
This simplifies into
$\det\left[\mathbf{R}\left(1;\lambda_{1},...,\lambda_{m}\right)\right]$
$\quad=\left(-1\right)^{n_{1}+...+n_{m}-mM}\left(\left(-1\right)^{n_{1}-N+1}\lambda_{1}+\lambda_{n_{1}}^{\left(N,M\right)}\right)...\left(\left(-1\right)^{n_{m}-N+1}\lambda_{m}+\lambda_{n_{m}}^{\left(N,M\right)}\right)$
×$\qquad\times\frac{b_{n_{1}}^{\left(N,M\right)}...b_{n_{m}}^{\left(N,M\right)}}{2^{m\left(m-1\right)/2}\left(M-1\right)^{m-1}\left(M-2\right)^{m-2}...(M-m+1)}D(n_{1},...,n_{m}),$ (6.33)
with $D(n_{1},...,n_{m})$ given by Eq. (6.27).
Comparing Eq. (6.33) with Eq. (6.21), condition (6.11) becomes
$\mathrm{sign}\left(\lambda_{1}...\lambda_{m}\right)=(-1)^{m\left(m-1\right)/2}\prod_{j=1}^{m}\mathrm{sign}\left((-1)^{n_{j}-N-M+1}\lambda_{j}+(-1)^{M}\lambda_{n_{j}}^{(N,M)}\right).$ (6.34)
As we have assumed that the ($m-1$)-step chain $N_{m-1}=\left(-n_{1}-1,\ldots,-n_{{m-1}}-1\right)$ is regular at each level, the condition above then implies that
$\mathrm{sign}\left(\lambda_{1}...\lambda_{m-1}\right)=(-1)^{(m-1)\left(m-2\right)/2}\prod_{j=1}^{m-1}\mathrm{sign}\left((-1)^{n_{j}-N-M+1}\lambda_{j}+(-1)^{M}\lambda_{n_{j}}^{(N,M)}\right).$ (6.35)
so that the regularity condition for the $m$-step chain reduces to
$\mathrm{sign}\left(\lambda_{m}\right)=(-1)^{m-1}\mathrm{sign}\left(\left(-1\right)^{n_{m}-N-M+1}\lambda_{m}+\left(-1\right)^{M}\lambda_{n_{m}}^{\left(N,M\right)}\right).$ (6.36)
Consider first the case where $m$ is even, which leads to
$\mathrm{sign}\left(\lambda_{m}\right)=\mathrm{sign}\left((-1)^{M}\left(\left(-1\right)^{n_{m}-N}\lambda_{m}-\lambda_{n_{m}}^{\left(N,M\right)}\right)\right).$ (6.37)
We then obtain the following possibilities:
$\left\{\begin{array}[]{c}(M,n_{m}-N)\in\left(2\mathbb{N}\right)^{2}:\lambda_{m}>\lambda_{n_{m}}^{(N,M)}\text{ or }\lambda_{m}<0;\\[5.69046pt] (M,n_{m}-N)\in 2\mathbb{N}\times\left(2\mathbb{N}+1\right):-\lambda_{m}^{(N,M)}<\lambda_{m}<0;\\[5.69046pt] (M,n_{m}-N)\in\left(2\mathbb{N}+1\right)\times 2\mathbb{N}:0<\lambda_{m}<\lambda_{n_{m}}^{(N,M)};\\[5.69046pt] (M,n_{m}-N)\in\left(2\mathbb{N}+1\right)^{2}:\lambda_{m}<-\lambda_{n_{m}}^{(N,M)}\text{ or }\lambda_{m}>0.\end{array}\right.$ (6.38)
Consider next the case where $m$ is odd, which leads to
$\mathrm{sign}\left(\lambda_{m}\right)=\mathrm{sign}\left((-1)^{M+1}\left(\left(-1\right)^{n_{m}-N}\lambda_{m}-\lambda_{n_{m}}^{\left(N,M\right)}\right)\right).$ (6.39)
We then obtain the following possibilities:
$\left\{\begin{array}[]{c}(M,n_{m}-N)\in\left(2\mathbb{N}\right)^{2}:0<\lambda_{m}<\lambda_{n_{m}}^{(N,M)};\\[5.69046pt] (M,n_{m}-N)\in 2\mathbb{N}\times\left(2\mathbb{N}+1\right):\lambda_{m}<-\lambda_{n_{m}}^{(N,M)}\text{ or }\lambda_{m}>0;\\[5.69046pt] (M,n_{m}-N)\in\left(2\mathbb{N}+1\right)\times 2\mathbb{N}:\lambda_{m}>\lambda_{n_{m}}^{(N,M)}\text{ or }\lambda_{m}<0;\\[5.69046pt] (M,n_{m}-N)\in\left(2\mathbb{N}+1\right)^{2}:-\lambda_{n_{m}}^{(N,M)}<\lambda_{m}<0.\end{array}\right.$ (6.40)
Provided conditions (6.38) or (6.40) are satisfied, the result will still be valid at the $m$th step.
B Rationally-extended potential
If $\lambda_{1},\lambda_{2},\ldots,\lambda_{m}$ satisfy the above-mentioned conditions, the $m$-step chain generates a regular potential
$V^{(N_{m})}(x;N,M;\lambda_{1},\ldots,\lambda_{m})$
$\quad=V(x;N,M)-2\frac{d^{2}}{dx^{2}}\log W(\psi_{-n_{1}-1}(x;N,M;\lambda_{1}),\ldots,\psi_{-n_{m}-1}(x;N,M;\lambda_{m})\mid x),$ (6.41)
which turns out to be a rational extension of $V(x;N-m,M-m)$,
$V^{(N_{m})}(x;N,M;\lambda_{1},\ldots,\lambda_{m})$
$\quad=V(x;N-m,M-m)+E_{-m}(N,M)$
$\qquad-8(1-z^{2})\frac{d^{2}}{dz^{2}}\log\left(\det\left[\mathbf{R}\left(z;\lambda_{1},...,\lambda_{m}\right)\right]\right)$
$\qquad+8z\frac{d}{dz}\log\left(\det\left[\mathbf{R}(z;\lambda_{1},...,\lambda_{m})\right]\right),$ (6.42)
where the matrix elements of $\mathbf{R}(z;\lambda_{1},...,\lambda_{m})$ are given in Eq. (6.10).
The derivation of (6.42) proceeds as that of (5.32) in Sec. V by using the identities
$W(\psi_{-n_{1}-1}(x;N,M;\lambda_{1}),\psi_{-n_{2}-1}(x;N,M;\lambda_{2}),\ldots,\psi_{-n_{m}-1}(x;N,M;\lambda_{m})\mid x)$
$\quad=\psi_{-1}^{m}(x;N,M)\left(\frac{dz}{dx}\right)^{m(m-1)/2}\det\left[\mathbf{R}\left(z;\lambda_{1},...,\lambda_{m}\right)\right],$ (6.43)
$\psi_{-1}^{m}(x;N,M)\left(\frac{dz}{dx}\right)^{m(m-1)/2}\propto\prod_{i=0}^{m-1}\psi_{-1}(x;N-i,M-i),$ (6.44)
$\sum_{i=0}^{m-1}E_{-1}(N-i,M-i)=E_{-m}(N,M),$ (6.45)
as well as the shape invariance property (5.30) of the TDPT potential.
For $V^{\left(N_{m}\right)}\left(x;N,M;\lambda_{1},\ldots,\lambda_{m}\right)$ to be a confining potential on $]0,\pi/2[$, we need to impose that $V\left(x;N-m,M-m\right)$ has this property, which is achieved for $N,M>m+1/2$, thence $N,M\geq m+1$. For such $N,M$ values, we note that $V^{\left(N_{m}\right)}\left(x;N,M;\lambda_{1},\ldots,\lambda_{m}\right)$ is actually strongly repulsive in both 00 and $\pi/2$, since the singularities are there of the type $g/x^{2}$ ($g\geq 3/4$) and $g/\left(\pi/2-x\right)^{2}$ ($g\geq 3/4$), respectively. This means that at each extremity, only one basis solution is quadratically integrable [33].
The corresponding eigenstates are given by
$\psi^{(N_{m})}_{k}(x;N,M;\lambda_{1},\ldots,\lambda_{m})$
$\quad=\frac{W(\psi_{-n_{1}-1}(x;N,M;\lambda_{1}),\ldots,\psi_{-n_{m}-1}(x;N,M;\lambda_{m}),\psi_{k}(x;N,M)\mid x)}{W(\psi_{-n_{1}-1}(x;N,M;\lambda_{1}),\ldots,\psi_{-n_{m}-1}(x;N,M;\lambda_{m}\mid x)},$ (6.46)
for $k=0,1,2,\dots$, and
$\psi^{(N_{m})}_{-n_{i}-1}(x;N,M;\lambda_{1},\ldots,\lambda_{m})$
$\quad=\frac{W\left(\psi_{-n_{1}-1}(x;N,M;\lambda_{1}),\ldots,\check{\psi}_{-n_{i}-1}(x;N,M;\lambda_{i}),\ldots,\psi_{-n_{m}-1}(x;N,M;\lambda_{m})\mid x\right)}{W(\psi_{-n_{1}-1}(x;N,M;\lambda_{1}),\ldots,\psi_{-n_{m}-1}(x;N,M;\lambda_{m})\mid x)},$ (6.47)
for $i=1$, 2, …, $m$, with corresponding energies $E_{k}(N,M)$, $k=0$, 1, 2, … and $E_{-n_{i}-1}$, $i=1$, 2, …, $m$, respectively. Here, $\check{\psi}_{-n_{i}-1}(x;N,M;\lambda_{i})$ means that $\psi_{-n_{i}-1}(x;N,M;\lambda_{i})$ is excluded from the Wronskian.
Eq. (4.7) can be generalized as [35]
$\frac{d^{l}}{dz^{l}}\left((1-z)^{N}(1+z)^{M}P_{k}^{(N,M)}(z)\right)=\left(-2\right)^{l}(k+l)_{\underline{l}}(1-z)^{N-l}(1+z)^{M-l}P_{k+l}^{(N-l,M-l)}(z).$ (6.48)
Using (2.5), (6.9), and (6.48), Eq. (6.46) can be rewritten as
$\psi_{k}^{(N_{m})}(x;N,M;\lambda_{1},\ldots,\lambda_{m})$
$\quad=\psi_{-1}(x;N,M)\left(\frac{dz}{dx}\right)^{m}$
×$\qquad\times\frac{W(p_{n_{1}}^{(-N,-M)}(z;\lambda_{1}),\ldots,p_{n_{m}}^{(-N,-M)}(z;\lambda_{m}),(1-z)^{N}(1+z)^{M}P_{k}^{(N,M)}(z)\mid z)}{W\left(p_{n_{1}}^{(-N,-M)}(z;\lambda_{1}),p_{n_{2}}^{(-N,-M)}(z;\lambda_{2}),\ldots,p_{n_{m}}^{(-N,-M)}(z;\lambda_{m})\mid z\right)}$
$\quad=W(p_{n_{1}}^{(-N,-M)}(z;\lambda_{1}),\ldots,p_{n_{m}}^{(-N,-M)}(z;\lambda_{m}),(1-z)^{N}(1+z)^{M}P_{k}^{(N,M)}(z)\mid z)$
×$\qquad\times\frac{\psi_{-1}(x;N-m,M-m)}{\det\left[\mathbf{R}\left(z;\lambda_{1},...,\lambda_{m}\right)\right]},$ (6.49)
where the last Wronskian can be written as
$W(p_{n_{1}}^{(-N,-M)}(z;\lambda_{1}),\ldots,p_{n_{m}}^{(-N,-M)}(z;\lambda_{m}),(1-z)^{N}(1+z)^{M}P_{k}^{(N,M)}(z)\mid z)$
$\quad=(1-z)^{N-m}(1+z)^{M-m}Q_{k}^{(n_{1},n_{2},\ldots,n_{m})}(z;N,M;\lambda_{1},\ldots,\lambda_{m})$ (6.50)
with
$Q_{k}^{(n_{1},n_{2},\ldots,n_{m})}(z;N,M;\lambda_{1},\ldots,\lambda_{m})$
$=$ $\left|\begin{array}[]{cccc}\mathbf{R}_{1,1}\left(z;\lambda_{1},...,\lambda_{m}\right)&...&\mathbf{R}_{1,m}\left(z;\lambda_{1},...,\lambda_{m}\right)&(1-z^{2})^{m}P_{k}^{(N,M)}(z)\\ ...&&...&...\\ \mathbf{R}_{m,1}\left(z;\lambda_{1},...,\lambda_{m}\right)&...&\mathbf{R}_{m,m}\left(z;\lambda_{1},...,\lambda_{m}\right)&\left(-2\right)^{m-1}(k+m-1)_{\underline{m-1}}\\ &&&\times(1-z^{2})P_{k+m-1}^{(N-m+1,M-m+1)}(z)\\ \mathbf{R}_{m+1,1}\left(z;\lambda_{1},...,\lambda_{m}\right)&...&\mathbf{R}_{m+1,m}\left(z;\lambda_{1},...,\lambda_{m}\right)&\left(-2\right)^{m}(k+m)_{\underline{m}}P_{k+m}^{(N-m,M-m)}(z)\end{array}\right|.$
We then get
$\psi_{k}^{(N_{m})}(x;N,M;\lambda_{1},\ldots,\lambda_{m})$
$\quad=\frac{\psi_{-1}(x;N-m,M-m)(1-z)^{N-m}(1+z)^{M-m}}{\det\left[\mathbf{R}\left(z;\lambda_{1},...,\lambda_{m}\right)\right]}\times Q_{k}^{(n_{1},n_{2},\ldots,n_{m})}(z;N,M;\lambda_{1},\ldots,\lambda_{m})$
$\quad=\frac{\psi_{0}(x;N-m,M-m)}{\det\left[\mathbf{R}\left(z;\lambda_{1},...,\lambda_{m}\right)\right]}\times Q_{k}^{(n_{1},n_{2},\ldots,n_{m})}(z;N,M;\lambda_{1},\ldots,\lambda_{m}),$ (6.52)
for $k=0,1,2,\ldots$.
Moreover,
$\psi_{-n_{i}-1}^{(N_{m})}(x;N,M;\lambda_{1},\ldots,\lambda_{m})$
$=\frac{1}{\psi_{-1}(x;N,M)\left(\frac{dz}{dx}\right)^{m-1}}$
×$\times\frac{W(p_{n_{1}}^{(-N,-M)}(z;\lambda_{1}),\ldots,\check{p}_{n_{i}}^{(-N,-M)}(z;\lambda_{i}),\ldots,p_{n_{m}}^{(-N,-M)}(z;\lambda_{m})\mid z)}{W\left(p_{n_{1}}^{(-N,-M)}(z;\lambda_{1}),\ldots,p_{n_{m}}^{(-N,-M)}(z;\lambda_{m})\mid z\right)},$ (6.53)
where $\check{p}_{n_{i}}^{(-N,-M)}(z;\lambda_{i})$ means that $p_{n_{i}}^{(-N,-M)}(z;\lambda_{i})$ is excluded from the Wronskian. This leads to
$\psi^{(N_{m})}_{-n_{i}-1}(x;N,M;\lambda_{1},\ldots,\lambda_{m})$
$\quad=\frac{\psi_{0}(x;N-m,M-m)}{\det[\mathbf{R}(z;\lambda_{1},\ldots,\lambda_{m})]}\times Q^{(n_{1},n_{2},\ldots,n_{m})}_{-n_{i}-1}(z;N,M;\lambda_{1},\ldots,\lambda_{m}),$ (6.54)
where $Q^{(n_{1},n_{2};\ldots,n_{m})}_{-n_{i}-1}(x;N,M;\lambda_{1},\ldots,\lambda_{m})$ is obtained from the determinant of $\mathbf{R}(z;\lambda_{1},\ldots,\lambda_{m})$ by suppressing the $m^{\rm th}$ row and the $i^{\rm th}$ column, i.e., by a minor that we demote by $\det[\mathbf{R}(z;\lambda_{1},\ldots,\lambda_{m})]^{m,i}$,
$Q^{(n_{1},n_{2};\ldots,n_{m})}_{-n_{i}-1}(x;N,M;\lambda_{1},\ldots,\lambda_{m})=\det[\mathbf{R}(z;\lambda_{1},\ldots,\lambda_{m})]^{m,i}.$ (6.55)
Due to the orthogonality properties of the $\psi_{k}^{(N_{m})}$, the polynomials $Q_{k}^{(n_{1},n_{2},\ldots,n_{m})}(z;N,M;\lambda_{1},\ldots,\lambda_{m})$, $k=-n_{1}-1$, …, $-n_{m}-1,0,1,2,\ldots$, form a set of orthogonal polynomials on ]1,1[]-1,1[] - 1 , 1 [ with respect of the measure
$\mu_{n_{1}n_{2}\ldots n_{m}}^{(-N,-M)}(z;\lambda_{1},\lambda_{2},\ldots,\lambda_{m})$
$\quad=\frac{(1-z)^{N-m}(1+z)^{M-m}}{\left(\det\left[\mathbf{R}\left(z;\lambda_{1},...,\lambda_{m}\right)\right]\right)^{2}}.$ (6.56)
VII CONCLUSION
In this article, we have studied the multi-step version of a previous construction of the TDPT potential regular rational extensions obtained by one-step DT using seed functions associated with the para-Jacobi polynomials of Calogero and Yi. We have shown that the eigenstates of such $m$-step extensions are expressed in terms of novel families of EOPs, orthogonal on ]1,+1[]-1,+1[] - 1 , + 1 [ and depending not only on $m$ discrete parameters, but also on $m$ real continuous parameters $\lambda_{1}$, $\lambda_{2}$, …, $\lambda_{m}$. The sets of parameters are related by some restrictions coming from the TDPT rational extensions regularity conditions, which we study in detail.
Looking for possible relationships between our results and those obtained by using other approaches for generating EOPs depending on an arbitrary number of continuous parameters would be an interesting open question for future work.
ACKNOWLEDGMENTS
CQ was supported by the Fonds de la Recherche Scientifique-FNRS under Grant No. 4.45.10.08.
AUTHOR DECLARATIONS
Conflict of interest
The authors have no conflict to disclose.
Author Contributions
Y. Grandati: Conceptualization (equal); Formal analysis (equal); Investigation (equal); Methodology (equal); Writing - original draft (equal); Writing - review & editing (equal).
C. Quesne: Conceptualization (equal); Formal analysis (equal); Investigation (equal); Methodology (equal); Writing - original draft (equal); Writing - review & editing (equal).
DATA AVAILABILITY
Data sharing is not applicable to this article since no new data were created or analyzed in this study.